OpenStudy (anonymous):

PLEASE CHECK ANSWER ! MEDAL WILL BE GIVEN ! . Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12 (14, 12) (12, 14) There is no solution. There are an infinite number of solutions. ANSWER IS C !

OpenStudy (lovelyanna):

Ya, its 100% C.

OpenStudy (anonymous):

thanks @LovelyAnna can you help me check the rest of my answers ?

OpenStudy (lovelyanna):

Ok, I'll try.

OpenStudy (anonymous):

Solve, using the substitution method. 2x + 3y = 1 y = x – 8 (0, –8) (2, –6) (4, –4) (5, –3) ANSWER - A & thaaansk so much

OpenStudy (lovelyanna):

Give me a minute.

OpenStudy (anonymous):

sure

OpenStudy (lovelyanna):

Can you tell me which grade this is?

OpenStudy (anonymous):

9th

OpenStudy (lovelyanna):

Oh, ok. I'm trying to figure it out, I'm still in 8th.

OpenStudy (anonymous):

woah . your smarter than me . i dont get these kind of things but really thaanks (: & i just started this homeschooling thing so thaanks

OpenStudy (lovelyanna):

Thanks! I do home school too, and I'm trying to do graphs, sloped and intercepts, there just too hard for me.. And for the 2nd question, I'd say A or D, not sure

OpenStudy (anonymous):

ok thanks and i got one more

OpenStudy (lovelyanna):

Ok, go ahead. I'll try my best.

OpenStudy (anonymous):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y There is no solution. There are an infinite number of solutions. The solution is (1, 10) The solution is (28, 0) ANSWER-D

OpenStudy (lovelyanna):

Hmm.. complected.. let me think about it

OpenStudy (anonymous):

ok

OpenStudy (lovelyanna):

Umm.. I dont know :/

OpenStudy (anonymous):

Solve, using the substitution method. y = 4x + 6 5x – y = 6 There are an infinite number of solutions. There is no solution. (18, 12) (12, 54) ANSWER- A ? ? i dont know ? ?

OpenStudy (lovelyanna):

Hold on

OpenStudy (anonymous):

ok

OpenStudy (lovelyanna):

I'm sorry, I haven't taken this yet and I don't know..

OpenStudy (anonymous):

ok thanks alot for your help !! ill just ask it again hopefully someone will know it ..

OpenStudy (lovelyanna):

Ok, Gl!